OpenStudy (anonymous):

Determine the slope of the graph of x2 = ln(xy) at the point (1, e) Do I start with implicit differentiation? Or is there another way? I do not know what to do.

OpenStudy (xmissalycatx):

@hero please help :)

hero (hero):

Try isolating y first.

hero (hero):

Afterwards, find y'

OpenStudy (benlindquist):

so many people

OpenStudy (anonymous):

\[y=\frac{ 1 }{ x ^{3} }\] @Hero So far so good?

OpenStudy (solomonzelman):

You can isolate the y, and as I did that I obtained \(\large\color{#000000 }{ \displaystyle y(x)=e^{x^2}/x }\) (Note, if you want to integrate that, then top and bottom •x, and then set u=x\(^2\),)

OpenStudy (solomonzelman):

9(No, you don't need to integrate this, this is not about integration at all, but just in case you wondered about it))

OpenStudy (solomonzelman):

If you want, you could do implicit differentiation as well. \(\large\color{#000000 }{ \displaystyle x^2 = \ln(xy) }\) \(\large\color{#000000 }{ \displaystyle x^2 = \ln(x)+\ln(y) }\) then, differentiate each component. (where the derivative of ln(y) deserves a y')

OpenStudy (anonymous):

Thanks!

OpenStudy (solomonzelman):

Just in case, (to nail the concept if it's not nailed already). \(\large\color{#000000 }{ \displaystyle \frac{ d}{dx}\left[~\ln f(x)~\right]=\frac{1}{f(x)}\times f'(x)=\frac{f'(x)}{f(x)} }\) (This is when you take an "ln" of a function; via the chain rule) And that above is exactly the equivalent of \(\large\color{#000000 }{ \displaystyle \frac{ d}{dx}\left[~\ln y~\right]=\frac{1}{y}\times y'=\frac{y'}{y} }\) Because y is a function of x (just as f(x) is), and therefore, just as f(x) it desrves a chain rule. ((In case you might have been wondering where this y' always comes from; hope I addressed this question))